• This week Alice played laser tag twice and climbed the rock wall 3 times, spending $31.
  • Last week Alice played laser tag 4 times and climbed the rock wall once, spending $27.
1
Expert's answer
2021-02-24T07:03:28-0500

Answer:

1. $7 for climbing the rock

Step-by-step explanation:

x is the number of laser tags

y is the number of times climbed the rock

a. 2x + 3y = 31

b. 4x + y =27

lets find y

Step 1

4x + y = 27

y = 27 -4x

step 2

replace y with 27 - 4x in the a.

2x + 3( 27 - 4x) = 31

2x + 81 - 12x = 31

2x -12x = 31 -81

x = 5

y = 27 - 4(5)

y = 7